# CMMCoT: Enhancing Complex Multi-Image Comprehension via Multi-Modal Chain-of-Thought and Memory Augmentation

### v1

@@ Abstract
While previous multimodal slow-thinking methods have demonstrated remarkable success in single-image understanding scenarios, their effectiveness becomes fundamentally constrained when extended to more complex multi-image comprehension tasks. This limitation stems from their predominant reliance on text-based intermediate reasoning processes. While for human, when engaging in sophisticated multi-image analysis, they typically perform two complementary cognitive operations: (1) continuous cross-image visual comparison through region-of-interest matching, and (2) dynamic memorization of critical visual concepts throughout the reasoning chain. Motivated by these observations, we propose the Complex Multi-Modal Chain-of-Thought (CMMCoT) framework, a multi-step reasoning framework that mimics human-like ”slow thinking” for multi-image understanding. Our approach incorporates two key innovations: (1) The construction of interleaved multimodal multi-step reasoning chains, which utilize critical visual region tokens, extracted from intermediate reasoning steps, as supervisory signals. This mechanism not only facilitates comprehensive cross-modal understanding but also enhances model interpretability. (2) The introduction of a test-time memory augmentation module that expands the model’s reasoning capacity during inference while preserving parameter efficiency. Furthermore, to facilitate research in this direction, we have curated a novel multi-image slow-thinking dataset. Extensive experiments demonstrate the effectiveness of our model. Code and dataset will be coming soon.
Machine Learning, ICML
1 Introduction
@@
Recent years have witnessed the rapid development of multi-modal large language models (MLLMs) in broad applications (Liu et al., 2024b; Wang et al., 2023, 2024a), i.e., multimodal recognition, localization, single and multi-image understanding. However, current MLLM methods employing the ”direct prediction” paradigm to generate answers end-to-end exhibit two critical limitations when confronted with intricate scenarios: (a) They systematically overlook latent evidential features embedded within cross-modal data patterns, resulting in erroneous predictions—even in state-of-the-art models like GPT-4V—when tackling intricate multimodal tasks(Yang et al., 2023), and (b) Their decision-making processes suffer from a pronounced lack of interpretability, undermining their reliability and transparency in critical applications.
Figure 1: CMMCot can enhance the model’s performance in handling complex multi-image tasks and leverage entity images to extract multi-image features during inference. Here is a case of complex multi-image scenarios.
@@
The remarkable success of the O1 model has catalyzed growing research interest in chain-of-thought (CoT) reasoning and ”slow thinking” mechanisms (Snell et al., 2024; Prystawski et al., 2024). These mechanisms yield substantial performance enhancements over conventional LLMs and significantly enhance mathematical reasoning, logical deduction capabilities, and sophisticated problem-solving capacities. However, unimodal slow-thinking methodologies exhibit inherent constraints when directly adapted to multimodal domains, primarily due to insufficient cross-modal spatial reasoning capacities for complex scene comprehension.
@@
Existing multimodal CoT approaches predominantly focus on employing external tools to annotate complete textual reasoning chains (Ni et al., 2024; Xu et al., 2024; Shao et al., 2024) or leveraging textual reasoning capabilities to guide visual inference (Du et al., 2025) while neglecting supervision of visual reasoning processes during training.
@@
Human cognition in multimodal contexts, particularly when processing multi-image compositions, operates through a dual-process cognitive mechanism: (1) parallel processing of linguistic semantics and visual signal interpretation, and (2) synergistic integration of textual deductive reasoning with active visual pattern mining. This cognitive architecture enables the iterative establishment of object-relation correspondences across cross-modal representations, ultimately synthesizing a coherent cross-modal reasoning framework through dynamic interaction between linguistic parsing and visual grounding.
Figure 2: The radar chart compares the performance of top MLLMs across six challenging multimodal benchmarks.
To mimic the aforementioned process, MVoT (Li et al., 2025) generates image tokens associated with the chain of thought. However, generating lengthy image token sequences inevitably leads to error accumulation, resulting in critical visual information decoding errors that compromise the model’s reasoning capabilities. While VoCoT (Li et al., 2024d) annotates object-centered multimodal reasoning paths, its direct extraction of token sequences from encoded image sequences based on coordinates inevitably introduces noise interference from features surrounding key objects. Moreover, the aforementioned datasets and models primarily focus on enhancing the slow reasoning capabilities of models in single-image scenarios, while leaving the more complex multi-image scenarios largely unexplored.
@@
In this paper, we address the challenge of slow thinking in complex multi-image scenarios. However, constructing multi-step reasoning paths for multi-image scenes presents two significant challenges: Complexity in Cross-image Visual Concept Tracking: Unlike single-image scenarios, comprehending multi-image scenes (Jiang et al., 2024) necessitates the correlation of visual objects across disparate images and the integration of textual information to facilitate comprehensive reasoning. Enhancement of Model’s Inference Capabilities during Testing: While test-time scaling has shown promise in boosting model reasoning (Ni et al., 2024; Xu et al., 2024) without parameter increases, its effectiveness diminishes in complex multi-image scenarios. Furthermore, simply scaling pre-trained model parameters is reaching a performance ceiling. Thus, exploring alternative methods to enhance model capabilities during testing for multi-image understanding is crucial.
@@
Building upon prior works (Brown et al., 2024; Wang et al., 2024b), we further propose the Retrieval-based Image Feature Reasoning Enhancement Module (RIFREM) module to enable comprehensive cross-modal deliberation during testing. Specifically, this module stores the Key and Value pairs of multi-image input tokens obtained after each decoder layer in a dedicated memory bank. When decoding specific coordinates, the model retrieves corresponding visual tokens using image IDs and coordinates, then computes attention between the query vectors (derived from subgraph token sequences processed through each decoder layer) and the stored multi-image Key/Value pairs in the memory bank.
@@
Furthermore, considering that existing works lack datasets meeting our requirements, we constructed a new dataset named CMMCoT-260k, which constitutes an innovative benchmark specifically designed for complex multi-image multimodal tasks, comprising 260,000 meticulously curated data instances. Distinct from conventional datasets, CMMCoT-260k’s uniqueness resides in its incorporation of explicit reasoning chains within each instance. These structured reasoning pathways not only facilitate deep semantic parsing of hybrid text-image data but also integrate spatial coordinates and entity-specific imagery, thereby enabling hierarchical reasoning analysis. Extensive experiments conducted on both multi-image and single-image benchmarks demonstrate the effectiveness of our model.
2 Dataset Construction
@@
Building upon multiple datasets GRIT (Peng et al., 2023), Flickr30k-Entities (Plummer et al., 2015), VoCoT (Li et al., 2024d), and MANTIS (Jiang et al., 2024), we have constructed a complex multi-image, multi-modal Chain-of-Thought (CoT) dataset, referred to as the CMMCoT-260K dataset, consisting of 260,000 data instances. This dataset encompasses four distinct types of data: Caption, Co-reference, Comparison, and Reason, each serving its unique purpose.
  • Caption: The caption data type is employed to generate natural language descriptions for images, serving as a direct expression of the semantic information contained within the image. It provides the model with the ability to comprehend and semantically parse image content, facilitating tasks such as annotating and generating descriptions for multiple images.
  • Co-reference: Co-reference data involves identifying and linking elements that refer to the same entity across different images. This data type assists the model in understanding entity consistency within a single image or among multiple images. It enhances the model’s ability to maintain consistent entity reasoning across multiple images, thereby strengthening its understanding of complex scenes.
  • Comparison: The comparison data type is utilized for conducting comparative analyses between multiple images, identifying similarities and differences among elements in different images. It improves the model’s comparative and contrastive analytical capabilities, enabling it to make more nuanced and accurate judgments in tasks involving multiple images, such as in visual question answering.
  • Reason: The reason data type pertains to more complex reasoning tasks, requiring the model to perform logical reasoning and decision-making based on relationships among multiple images. This typically necessitates contextual understanding and causal relationship analysis between images. It enhances the model’s capacity to solve advanced reasoning problems, making it applicable for understanding more complex task contexts.
For the Caption and Co-reference tasks, we developed a straightforward pipeline that integrates data annotated with scene information from individual images, thereby forming our CMMCoT-260K dataset - an innovative collection of 260K data instances specifically designed for complex multi-image multimodal tasks. Unlike conventional datasets, CMMCoT-260K distinguishes itself through its structured reasoning chains that accompany each instance. These reasoning chains provide deep contextual analysis for mixed text-image data while incorporating critical spatial coordinates and entity-level visual grounding, enabling comprehensive multi-level reasoning capabilities. The dataset architecture not only enhances semantic understanding of visual content but also ensures cross-image entity consistency and complex scene interpretation.
In contrast, for more challenging multi-image tasks such as Comparison and Reasoning, we designed a more complex data processing pipeline. This process includes several steps: ”constructing the QA rationale chain (Zelikman et al., 2024); entity extraction; entity detection; entity relationship summarization.” To ensure robust multimodal analysis, we systematically curated and extended high-quality resources from multiple established datasets including GRIT, Flickr30k-Entities, VoCoT, and Mantis, ultimately forming four core data categories. The complete implementation of this pipeline helps us accurately capture and understand the complex relationships between different images.
Specifically, on the ImageCoDe dataset, we innovatively proposed an ”N-in-One” synthesis method. This technique involves combining a fixed number of images into a single image, which not only simplifies the processing and construction of multi-image data but also allows us to naturally associate image indices with their corresponding scene information.
@@
The CMMCoT-260K dataset significantly advances multimodal research by providing both comprehensive data support and expanded application scenarios for developing novel algorithms. In summary, through these methods and technical approaches, we have successfully transformed multi-image QA datasets into a comprehensive and powerful multi-image multimodal CMMCoT dataset, establishing an essential foundation for sophisticated multimodal reasoning systems.
@@
3 Methods
Baseline: Along with the proposed dataset, we also develop a novel multi-modal framework named CMMCoT, which employs qwen2-vl (Wang et al., 2024a) as the baseline model. The primary reason for selecting Qwen2-VL is its trained Vision Transformer (ViT) that supports Naive Dynamic Resolution, enabling the processing of images at arbitrary resolutions and dynamically converting them into a variable number of visual tokens. Furthermore, Qwen2-VL introduces Multimodal Rotary Position Embedding (M-RoPE), which effectively models the positional information of multimodal inputs. This approach reduces the positional ID values of images and videos, allowing the model to extrapolate to longer sequences during inference.
Multimodal Sequence Representation: CMMCoT represents complex multi-image and textual data in an interleaved visual-textual format. For complex multi-image tasks, it is essential to compare and contrast the associations and differences among different images, between different entities within the same image, and between different entities across different images during the reasoning process. Therefore, we introduce special image index tokens to refer to specific input images, formatted as <IMG>0</IMG>. Here, <IMG> and </IMG> are special tokens indicating the start and end of an image index. This indexing logic can be extended to accommodate more images.
Additionally, we represent different entities using coordinates and visual tokens. The coordinate format is similar to <|box_start|>(<x_min>, <y_min>) , (<x_max>, <y_max>)<|box_end|>, where <|box_start|> and <|box_end|> are special markers indicating the beginning and end of coordinate information. We use bounding boxes (<x_min>, <y_min>) and (<x_max>, <y_max>) as the entity coordinates, with x and y normalized relative to the image dimensions, ranging from 0 to 1000. The coordinate values are tokenized and embedded as textual data.
For visual information pertaining to multiple images and entities, we use special markers <|vision_start|> and <|vision_end|> to denote the start and end positions of visual content. During the training process, we obtain entity images based on their coordinates and image indices and encode each entity image using a visual encoder. When encoding the entity images through the visual encoder, we limit the minimum resolution of the entities to 512 pixels, which allows us to extract more detailed features from the entity images (Li et al., 2024c).
@@ 3 Methods
RIFREM: Previous studies (Snell et al., 2024; Prystawski et al., 2024) have demonstrated that expanding model computational capacity during inference can significantly enhance performance metrics. For multi-image tasks, a natural approach involves comparing critical visual regions with relevant areas in other images through joint reasoning when identifying pivotal visual attention zones.
@@
=$\mathcal{M}=\{(K_{l}^{(i)},V_{l}^{(i)})\}_{l=1,2,...,L;i=1,2,...,N}$ (1)
where $K_{l}^{(i)},V_{l}^{(i)}$ represent the keys and values from the $l$-th decoder layer of the $i$-th multi-image sequence, $L$ is the total number of decoder layers, and $N$ is the number of multi-image sequences processed.
During CoT prediction, when encountering the </IMG> token in the reasoning process, we inject entity image tokens into decoder layers and extract their query vectors based on position IDs. These queries then engage in cross-attention computation with corresponding key-value pairs retrieved from the memory bank. The attention mechanism can be formally described as scaled dot-product attention:
@@
Table 1: Performance Comparison of SOTA Models across Multi-image Benchmarks
Models Parameters BLINK Mantis NLVR2 MVBench Q-Bench Avg
LLava-v1.5 7B 37.1 41.9 52.1 36.0 53.9 44.2
LLava-v1.6 7B 39.6 45.6 58.9 40.9 58.9 48.8
LLava-v1.5-MIA-DPO 7B 42.9 44.2 54.2 39.5 - -
Mantis-Idefics2 8B 49.1 57.1 89.7 51.4 75.3 64.5
llava-ov 7B 48.2 64.2 89.4 56.7 74.5 66.6
qwen2-vl 7B 51.8 64.8 84.0 50.5 71.2 64.5
Ours 7B 52.3 70.9 88.7 51.3 72.2 67.1
qwen2-vl 2B 43.6 37.7 74.1 37.5 58.5 50.3
Ours 2B 44.8 39.8 76.9 40.1 61.7 52.7
Table 2: Performance Comparison of SOTA Models across Single-image Benchmarks
Models Param. MMMU MMStar SQA RealWorldQA MME POPE HallBench Avg
@@
Ours 7B 48.2 58.9 88.9 65.0 81.5 86.3 60.5 69.9
@@
Ours 2B 39.4 49.2 78.6 62.6 65.7 84.0 52.5 61.7
4 Experiments
4.1 Evaluation Benchmarks
NLVR2 (Suhr et al., 2018) evaluates visual-textual reasoning through cross-image logical analysis, presenting paired images with declarative statements for propositional verification. It employs constrained selection on the standard test-public partition.
Qbench (Wu et al., 2023) assesses multimodal models’ critical analysis of benchmarking parameters, focusing on visual perception in quality assessment tasks. Our experiment employs its Qbench2-A1-dev subset with multiple-choice comparative visual analysis.
Mantis-Eval contains 217 multi-image analysis tasks across conceptual domains (dimensional assessment, mass estimation). Manually curated from web-sourced images, it combines closed and open-response formats evaluated on standardized test partitions.
BLINK (Fu et al., 2025) tests rapid visual cognition (depth relationships, feature matching, digital forensics, spatial-temporal reasoning). It includes multi-image perceptual similarity tasks, measured through standardized validation protocols.
MVBench (Li et al., 2024b) evaluates temporal reasoning in video comprehension through 20 tasks analyzing frame dynamics. Using 8-frame sampling per video, results are quantified via standardized test-set measurements.
Single Image Benchmarks: We also conduct experiments on single-image tasks to further demonstrate the effectiveness of our proposed model: MMMU (Yue et al., 2024), MMStar (Chen et al., 2024), SQA (Lu et al., 2022), RealWorldQA, MME (Yin et al., 2023), POPE (Li et al., 2023), HallBench (Guan et al., 2023).
4.2 Experiments Setting
We conduct supervised fine-tuning (SFT) based on the Qwen2-VL-7B architecture. The optimization process employs AdamW with $\beta=0.95$ and a weight decay of 0.1, coupled with a cosine learning rate scheduler. The training framework utilizes the DeepSpeed ZeRO-3 strategy for efficient parameter optimization.
@@ 4.2 Experiments Setting
The training regimen consists of two distinct phases: Stage 1: Initial learning rate of 1e-5 with 2 training epochs. Stage 2: Reduced learning rate of 1e-6 with 1 training epoch at batch size 256. This two-phase approach enables progressive refinement of model parameters, where the first stage establishes coarse-grained feature representations and the second stage performs fine-grained adjustment with reduced stochasticity through smaller learning rates and larger batch sizes.
4.3 Compared Methods
@@
To evaluate and compare the effectiveness of multi-image models, we conducted comprehensive benchmarking on several state-of-the-art architectures: LLaVA-v1.5 (Liu et al., 2024a), LLaVA-v1.6 (Li et al., 2024a), LLaVA-v1.5-MIA-DPO (Liu et al., 2024c), Mantis-Idefics2 (Jiang et al., 2024), and LLaVA-OV (Li et al., 2024a). Our evaluation framework simultaneously addresses both multi-image and single-image tasks, enabling comparative analysis of model performance across different input configurations.
4.4 Performance Evaluation

### v2

@@ Abstract
While previous multimodal slow-thinking methods have demonstrated remarkable success in single-image understanding scenarios, their effectiveness becomes fundamentally constrained when extended to more complex multi-image comprehension tasks. This limitation stems from their predominant reliance on text-based intermediate reasoning processes. While humans, when engaging in sophisticated multi-image analysis, typically perform two complementary cognitive operations: (1) continuous cross-image visual comparison through region-of-interest matching, and (2) dynamic memorization of critical visual concepts throughout the reasoning chain. Motivated by these observations, we propose the Complex Multi-Modal Chain-of-Thought (CMMCoT) framework, a multi-step reasoning framework that mimics human-like ”slow thinking” for multi-image understanding. Our approach incorporates two key innovations: (1) The construction of interleaved multimodal multi-step reasoning chains, which utilize critical visual region tokens, extracted from intermediate reasoning steps, as supervisory signals. This mechanism not only facilitates comprehensive cross-modal understanding but also enhances model interpretability. (2) The introduction of a test-time memory augmentation module that expands the model’s reasoning capacity during inference while preserving parameter efficiency. Furthermore, to facilitate research in this direction, we have curated a novel multi-image slow-thinking dataset. Extensive experiments demonstrate the effectiveness of our model. Code is available at https://github.com/zhangguanghao523/CMMCoT.
Introduction
@@
Recent years have witnessed the rapid advancement of generative (he2025mars; esser2024scalingrectifiedflowtransformers; wang2025mint; liu2025tfcustom; ma2024followyouremoji; ma2025followcreation; feng2025dit4edit) and multimodal understanding models (wang2025mint; Wang2024Qwen2VLEV). In particular, multi-modal large language models (MLLMs) have achieved remarkable breakthroughs across various multimodal tasks (liu2024visual; wang2023cogvlm; wang2024qwen2), such as multimodal recognition, localization, and single- as well as multi-image understanding. However, current MLLM methods employing the ”direct prediction” paradigm to generate answers end-to-end exhibit two critical limitations when confronted with intricate scenarios: (a) They systematically overlook latent evidential features embedded within cross-modal data patterns, resulting in erroneous predictions—even in state-of-the-art models like GPT-4V(yang2023set), and (b) Their decision-making processes suffer from a pronounced lack of interpretability, undermining their reliability and transparency in critical applications.
@@
The remarkable success of the O1 model has catalyzed growing research interest in chain-of-thought (CoT) reasoning and ”slow thinking” mechanisms (snell2024scaling; prystawski2024think). These mechanisms yield substantial performance enhancements over conventional LLMs and significantly enhance mathematical reasoning and logical deduction capabilities. However, unimodal slow-thinking methodologies exhibit inherent constraints when directly adapted to multimodal domains, primarily due to insufficient cross-modal spatial reasoning capacities for complex scene comprehension.
@@
Existing multimodal CoT approaches predominantly focus on employing external tools to annotate complete textual reasoning chains (ni2024visual; xu2024llava; shao2024visual) or leveraging textual reasoning capabilities to guide visual inference (du2025virgo) while neglecting supervision of visual reasoning processes during training.
@@
Human cognition in multimodal contexts, particularly when processing multi-image compositions, operates through a dual-process cognitive mechanism: (1) parallel processing of linguistic semantics and visual signal interpretation, and (2) synergistic integration of textual deductive reasoning with active visual pattern mining. This cognitive architecture enables the iterative establishment of object-relation correspondences across cross-modal representations, ultimately synthesizing a coherent cross-modal reasoning framework through dynamic interaction between linguistic parsing and visual grounding. Although previous works like MVoT (li2025imagine) and VoCoT (li2024vocot) have been proposed to mimic the aforementioned process, these methods primarily focus on enhancing the slow reasoning capabilities of models in single-image scenarios, while leaving the more complex multi-image scenarios largely unexplored.
Figure 1: Data generation process of our proposed CMMCoT dataset. The construction of the CMMCoT dataset mainly concerns four parts: the generation of QA rationale chains, the extraction of textual entities, the detection and validation of visual entities, and the spatial fusion and summarization of entity groundings.
@@
Thus in this paper, we address the challenge of slow thinking in complex multi-image scenarios. However, constructing multi-step reasoning paths for multi-image scenes presents two significant challenges: Complexity in Cross-image Visual Concept Tracking: Unlike single-image scenarios, comprehending multi-image scenes (jiang2024mantis) necessitates the correlation of visual objects across disparate images and the integration of textual information to facilitate comprehensive reasoning. Enhancement of Model’s Inference Capabilities during Testing: While test-time scaling has shown promise in boosting model reasoning (ni2024visual; xu2024llava) without parameter increases, its effectiveness diminishes in complex multi-image scenarios. Furthermore, simply scaling pre-trained model parameters is reaching a performance ceiling. Thus, exploring alternative methods to enhance model capabilities during testing for multi-image understanding is crucial.
@@
Building upon prior works (brown2024large; wang2024augmenting), we further propose the Retrieval-based Image Feature Reasoning Enhancement Module (RIFREM) module to enable comprehensive cross-modal deliberation during testing. Specifically, this module stores the Key and Value pairs of multi-image input tokens obtained after each decoder layer in a dedicated memory bank. When decoding specific coordinates, the model retrieves corresponding visual tokens using image IDs and coordinates, then computes attention between the query vectors (derived from subgraph token sequences processed through each decoder layer) and the stored multi-image Key/Value pairs in the memory bank.
@@
Furthermore, considering that existing works lack datasets meeting our requirements, we constructed a new dataset named CMMCoT-260k, which constitutes an innovative benchmark specifically designed for complex multi-image multimodal tasks, comprising 260,000 meticulously curated data instances. Distinct from conventional datasets, CMMCoT-260k’s uniqueness resides in its incorporation of explicit reasoning chains within each instance. These structured reasoning pathways not only facilitate deep semantic parsing of hybrid text-image data but also integrate spatial coordinates and entity-specific imagery, thereby enabling hierarchical reasoning analysis.
Extensive experiments conducted on both multi-image and single-image benchmarks demonstrate the effectiveness of our model. Moreover, visualization experiments have also demonstrated that compared to the traditional GPT-4o and Qwen2.5-VL, our model not only improves the accuracy of responses but also significantly enhances the interpretability of the intermediate reasoning process, making it more accessible for human understanding.
Dataset Construction
@@
Building upon multiple datasets including GRIT (peng2023kosmos), Flickr30k-Entities (plummer2015flickr30k), VoCoT (li2024vocot), and MANTIS (jiang2024mantis), we have constructed a complex multi-image, multi-modal Chain-of-Thought (CoT) dataset, referred to as the CMMCoT-260K dataset. It consists of 260,000 instances, encompassing four distinct task types: Caption, Co-reference, Comparison, and Reason.
For the Caption and Co-reference tasks, we employed a straightforward data integration pipeline. In contrast, for the more challenging Comparison and Reason tasks, we designed a more complex data processing pipeline that includes the following steps:
  • Constructing QA Rationale Chains: We generate CoT annotations through a fully automated pipeline based on the methodology from (zelikman2024star). Initially, GPT-4o is used to generate a preliminary reasoning chain and answer from the question. If the answer is correct, the chain is retained. If incorrect, the question is paired with the pre-annotated correct answer and re-fed into GPT-4o to generate a refined rationale. This process uses answer correctness as a quality filter to ensure the reliability of the generated rationales.
  • Entity Extraction: The Qwen3-235B-A22B model is utilized to extract textual entities from QA dialogues, providing a foundation for subsequent entity localization and relationship summarization.
  • Entity Detection: To ensure localization accuracy, we adopt a two-stage validation mechanism. First, the Qwen-VL-max model generates initial bounding boxes for entities. Subsequently, GPT-4o calculates the Intersection over Union (IoU) for these boxes, and only samples with an IoU $\geq$ 0.9 are retained to ensure precision.
  • Entity Relationship Summarization: When a single entity corresponds to multiple bounding boxes across images or detections, we apply spatial fusion. The smallest top-left and largest bottom-right coordinates from all relevant boxes are used to create a single, unified bounding box. This approach simplifies the CoT structure while maintaining complete entity coverage, enhancing the clarity and logical coherence of the reasoning chain.
@@
In summary, through this pipeline, we have successfully transformed multi-image QA datasets into a comprehensive and powerful multi-image, multi-modal CMMCoT dataset, establishing an essential foundation for sophisticated multi-modal reasoning systems.
@@
Methods
Baseline
Along with the proposed dataset, we also develop a novel multi-modal framework named CMMCoT, which employs Qwen2-VL (wang2024qwen2) as the baseline model. The primary reason for selecting Qwen2-VL is its trained Vision Transformer (ViT) that supports Naive Dynamic Resolution, enabling the processing of images at arbitrary resolutions and dynamically converting them into a variable number of visual tokens. Furthermore, Qwen2-VL introduces Multimodal Rotary Position Embedding (M-RoPE), which effectively models the positional information of multimodal inputs. This approach reduces the positional ID values of images and videos, allowing the model to extrapolate to longer sequences during inference.
Multimodal Sequence Representation
CMMCoT represents complex multi-image and textual data in an interleaved visual-textual format. For complex multi-image tasks, it is essential to compare and contrast the associations and differences among different images, between different entities within the same image, and between different entities across different images during the reasoning process. Therefore, we introduce special image index tokens to refer to specific input images, formatted as <IMG>0</IMG>. Here, <IMG> and </IMG> are special tokens indicating the start and end of an image index. This indexing logic can be extended to accommodate more images.
Additionally, we represent different entities using coordinates and visual tokens. The coordinate format is similar to <|box_start|>($x_{0}$>, $y_{0}$>) , ($x_{1}$>, $y_{1}$>)<|box_end|>, where <|box_start|> and <|box_end|> are special markers indicating the beginning and end of coordinate information. We use bounding boxes ($x_{0}$>, $y_{0}$>) and ($x_{1}$>, $y_{1}$>) as the entity coordinates, with x and y normalized relative to the image dimensions, ranging from 0 to 1000. The coordinate values are tokenized and embedded as textual data.
For visual information pertaining to multiple images and entities, we use special markers <|vision_start|> and <|vision_end|> to denote the start and end positions of visual content. During the training process, we obtain entity images based on their coordinates and image indices and encode each entity image using a visual encoder. When encoding the entity images through the visual encoder, we limit the minimum resolution of the entities to 512 pixels, which allows us to extract more detailed features from the entity images (li2024mini).
@@ Methods
Model Params BLINK Mantis NLVR2 MVBench Q-Bench Avg
LLaVA-v1.5 7B 37.1 41.9 52.1 36.0 53.9 44.2
LLaVA-v1.6 7B 39.6 45.6 58.9 40.9 58.9 48.8
LLaVA-v1.5-MIA-DPO 7B 42.9 44.2 54.2 39.5
LLaVA-OV 7B 48.2 64.2 89.4 56.7 74.5 66.6
Mantis-Idefics2 8B 49.1 57.1 89.7 51.4 75.3 64.5
InternVL3 8B 55.5 70.1 88.5 75.4 75.9 73.1
Qwen2-VL 7B 51.8 64.8 84.0 50.5 71.2 64.5
Qwen2-VL (Ours) 7B 52.3 70.9 88.7 51.3 72.2 67.1
Qwen2.5-VL 7B 55.3 69.8 88.3 74.7 77.7 73.2
Qwen2.5-VL (Ours) 7B 56.8 72.2 89.9 75.8 78.5 74.6
InternVL3 2B 50.3 65.9 85.4 70.4 71.5 68.7
Qwen2-VL 2B 43.6 37.7 74.1 37.5 58.5 50.3
Qwen2-VL (Ours) 2B 44.8 39.8 76.9 40.1 61.7 52.7
Qwen2.5-VL 3B 49.1 62.7 86.2 71.3 74.9 68.8
Qwen2.5-VL (Ours) 3B 51.4 68.5 88.9 73.1 75.2 71.4
Table 1: Performance comparison of SOTA models on multi-image benchmarks. Rows in light gray background and darker gray background represent Qwen2-VL and Qwen2.5-VL variants, respectively. Best results per benchmark are in bold.
RIFREM
Previous studies (snell2024scaling; prystawski2024think) have demonstrated that expanding model computational capacity during inference can significantly enhance performance metrics. For multi-image tasks, a natural approach involves comparing critical visual regions with relevant areas in other images through joint reasoning when identifying pivotal visual attention zones.
@@
$\mathcal{M}=\{(K_{l}^{(i)},V_{l}^{(i)})\}_{l=1,2,...,L;i=1,2,...,N}$ (1)
where $K_{l}^{(i)},V_{l}^{(i)}$ represents the keys and values from the $l$-th decoder layer of the $i$-th multi-image sequence, $l$ is the number of decoder layers, and $N$ is the number of multi-image sequences processed.
During CoT prediction, when encountering the </IMG> token in the reasoning process, we inject entity image tokens into decoder layers and extract their query vectors based on position IDs. These queries then engage in cross-attention with corresponding key-value pairs retrieved from the memory bank. The attention mechanism can be formally described as scaled dot-product attention:
@@
Models Params MMMU MMStar SQA RealWorldQA MME POPE HallBench Avg
@@
Qwen2-VL (Ours) 7B 48.2 58.9 88.9 65.0 81.5 86.3 60.5 69.9
Qwen2.5-VL 7B 58.6 63.9 89.0 68.4 82.6 85.9 51.9 71.4
Qwen2.5-VL (Ours) 7B 57.5 66.4 96.8 71.6 83.5 89.2 63.6 75.5
InternVL3 2B 48.6 61.1 95.8 65.1 78.1 90.1 41.9 68.7
@@
Qwen2-VL (Ours) 2B 39.4 49.2 78.6 62.6 65.7 84.0 52.5 61.7
Qwen2.5-VL 3B 53.1 55.8 81.4 65.5 78.6 85.9 46.6 66.7
Qwen2.5-VL (Ours) 3B 53.0 57.3 88.6 67.2 79.8 87.1 48.6 68.8
Table 2: Performance Comparison of SOTA Models across Single-image Benchmarks. The Qwen2-VL series are highlighted with light gray background and Qwen2.5-VL series with darker gray background, while other models use white background. Bold numbers indicate the best performance for each benchmark.
Grounding MCoT RIFREM BLINK Mantis NLVR2 MVBench Q-Bench Avg
55.3 69.8 88.3 74.7 77.7 73.2
55.2 70.4 88.7 74.5 77.9 73.3
57.1 71.6 89.4 75.4 78.7 74.4
56.8 72.2 89.9 75.8 78.5 74.6
Table 3: Performance (%) of Qwen2.5-VL-7B variants across benchmarks with different module combinations
Experiments
Evaluation Benchmarks
NLVR2 (suhr2018corpus) evaluates visual-textual reasoning through cross-image logical analysis, presenting paired images with declarative statements for propositional verification. It employs constrained selection on the standard test-public partition. Qbench (wu2023q) assesses multimodal models’ critical analysis of benchmarking parameters, focusing on visual perception in quality assessment tasks. Our experiment employs its Qbench2-A1-dev subset with multiple-choice comparative visual analysis. Mantis-Eval contains 217 multi-image analysis tasks across conceptual domains (dimensional assessment, mass estimation). Manually curated from web-sourced images, it combines closed and open-response formats evaluated on standardized test partitions. BLINK (fu2025blink) tests rapid visual cognition (depth relationships, feature matching, digital forensics, spatial-temporal reasoning). It includes multi-image perceptual similarity tasks, measured through standardized validation protocols. MVBench (li2024mvbench) evaluates temporal reasoning in video comprehension through 20 tasks analyzing frame dynamics. Using 8-frame sampling per video, results are quantified via standardized test-set measurements. Single Image Benchmarks: We also conduct experiments on single-image tasks to further demonstrate the effectiveness of our proposed model: MMMU (yue2024mmmu), MMStar (chen2024we), SQA (lu2022learn), RealWorldQA, MME (yin2023survey), POPE (li2023evaluating), HallBench (guan2023hallusionbench).
Experiments Setting
@@
We conduct supervised fine-tuning (SFT) based on the Qwen2.5-VL-7B architecture. The optimization process employs AdamW with $\beta=0.95$ and a weight decay of 0.1, coupled with a cosine learning rate scheduler. The training framework utilizes the DeepSpeed ZeRO-3 strategy for efficient parameter optimization. The training regimen consists of two distinct phases: Stage 1: Initial learning rate of 1e-5 with 2 training epochs. Stage 2: Reduced learning rate of 1e-6 with 1 training epoch at batch size 256. This two-phase approach enables progressive refinement of model parameters, where the first stage establishes coarse-grained feature representations and the second stage performs fine-grained adjustment with reduced stochasticity through smaller learning rates and larger batch sizes.
Compared Methods
@@
To evaluate and compare the effectiveness of multi-image models, we conducted comprehensive benchmarking on several state-of-the-art architectures: LLaVA-v1.5 (liu2024improved), LLaVA-v1.6 (li2024llava), LLaVA-v1.5-MIA-DPO (liu2024mia), Mantis-Idefics2 (jiang2024mantis), LLaVA-OV (li2024llava), Qwen2-VL (Wang2024Qwen2VLEV), InternVL3 (zhu2025internvl3) and Qwen2.5-VL (Bai2025Qwen25VLTR). Our evaluation framework simultaneously addresses both multi-image and single-image tasks, enabling comparative analysis of model performance across different input configurations.
Performance Evaluation